# Viologen-based supramolecular crystal gels: gelation kinetics and sensitivity to temperature

### v1

@@
Supramolecular crystal gels, a subset of molecular gels, form through self-assembly of low molecular weight gelators into interconnecting crystalline fibers, creating a three-dimensional soft solid network. This study focuses on the formation and properties of viologen-based supramolecular crystalline gels. It aims to answer key questions about the tunability of network properties and the origin of these properties through in-depth analyses of the gelation kinetics triggered by thermal quenching. Experimental investigations, including UV-Vis absorption spectroscopy, rheology, microscopy and scattering measurements, contribute to a comprehensive and self-consistent understanding of the system kinetics. We confirm that the viologen-based gelators crystallize into nanometer radius hollow tubes that assemble into micro to millimetric spherulites. We then show that the crystallization follows the Avrami theory and is based on pre-existing nuclei. We also establish that the growth is interface controlled leading to the hollow tubes to branch into spherulites with fractal structures. Finally, we demonstrate that the gel properties can be tuned depending on the quenching temperature. Lowering the temperature results in the formation of denser and smaller spherulites. In contrast, the gels elasticity is not significantly affected by the quench temperature, leading us to hypothesize that the spherulites densification occurs at the expense of the connectivity between spherulite.
@@ I Introduction
Figure 1: Characterization of the fibrous network in gel samples (a) Chemical structure of the viologen gelator $Chol_{2}V^{2+}$, composed of a 4,4’-bipyridinium unit bearing two cholesteryl substituants. (b) Microstructure of the gel assessed by small-angle X-ray scattering measurements. The scattering intensity $I(q)$ (circles) at high $q$ is best modeled by the form factor of a hollow cylinders (green curve). The inset figure represents the scattering length density profile of the hollow cylinder, with $\Delta\rho=\rho_{1}-\rho_{2}=5.10^{-7}\leavevmode\nobreak\ \rm{\AA}^{-2}$ the difference of scattering length densities between the solvent and the viologen gelator. $r_{i}=5.2\leavevmode\nobreak\ \rm nm$ and $t_{s}=2.6\leavevmode\nobreak\ \rm nm$ are the core radius and shell thickness of the cylinder, respectively. (c)-(d) Fibrous structure of the gel observed by transmission electron cryomicroscopy and atomic force microscopy, respectively.
@@
However, it remains unclear how these spherulites contribute to the formation of a 3D network. Do the branches entangle between neighboring spherulites, or do branches stick to one another due to attractive interactions? Predictions [46, 47, 48, 49] were made at the macroscopic level to establish a direct link between the crystal gel structure and its mechanical properties. In this context, the spherulites are best described as Cayley trees of size $\xi$ and fractal dimension $d_{f}$ composed of branching segment of length $l$ and radius $a$. Based on simulations, Shi et al. [43] assimilated the gel elasticity to that of the spherulites. In [46, 47, 48, 49], the authors proposed a model derived from the Avrami model [35] that links the temporal evolution of the elastic modulus to the fractal dimension $d_{f}$ of the crystal gel network. The model establishes that $X$, the ratio of the volume fraction of crystal material at time $t$ to the volume fraction at $t\to\infty$, is linked to the evolution of the norm of the complex viscosity $\eta^{*}$ through the equations $X=\frac{|\eta^{*}|-|\eta^{*}(0)|}{|\eta^{*}_{\infty}|-|\eta^{*}(0)|}$ and $\ln(1-X)=k(t-t_{g})^{d_{f}}$, assuming $n_{N}=0$ and $n_{G}=1$ et Eq. 2.
@@ II.1 Viologen-based gelator
The viologen-based gelator used for all experiments is a cholesteryl-substituted viologen noted $Chol_{2}V^{2+}$ (Fig. 1a). It was obtained as a bromide salt $V(Br)_{2}$ using a protocol described previously  [6]. The gelation ability of $Chol_{2}V^{2+}$ was first reported by Xue et al., who demonstrated that the gel fibrillar network could be efficiently used as a template for obtaining monodisperse mesoporous silica [28]. Our previous work showed that self-assembly of $Chol_{2}V^{2+}$ in 1-Pentanol relies on the formation of charge transfer (CT) complexes between $Chol_{2}V^{2+}$ (acceptor unit) and halide ions, $Br^{-}$ or $I^{-}$ (donor unit). The anions associated with the dicationic species $Chol_{2}V^{2+}$ play a major role in the self-assembly process and that the rheological, optical and electrical properties of the gels can be tuned by the addition of salts. We also showed that the addition of lithium bis(trifluoromethanesulfonyl) ($LiTFSI$) helps to increase the solubility of $Chol_{2}V^{2+}$ in pentanol but that too much TFSI anions can prevent gelation. Addition of $TBAI$ also increases the amount of donors species available to form the $[Chol_{2}V^{2+}/I^{-}]$ CT complexes promoting gelation but yields opaque gels displaying high gelation temperature. An optimal composition of $TBAI$ and $LiTFSI$ has thus been determined to obtain a translucent and robust gel featuring a moderate gelation temperature and a high ionic conductivity [6]. This study specifically focuses on this optimal composition consisting of $Chol_{2}V(Br)_{2}$ (1.3 mM, 0.2 wt%), $TBAI$ (50 mM), and $LiTFSI$ (100 mM) in 1-Pentanol. In all experiments, the sol-gel transition was induced by thermally quenching homogeneous solutions at a rate of $\sim 0.6^{\circ}$/s from 80 C to a final temperature $T$ using a Peltier module. The shallower quench where carried out to a final temperature $T=50^{\circ}$ while the deeper quenches where carried out down to $T=20^{\circ}$.
@@ II.4 Bright field microscopy
Microscopy measurements were performed using a Nikon Ti-eclipse inverted light microscope operating in bright-field mode. A Zyla camera from Andor, equipped with a 10× objective, was employed to capture images of the samples. A 455 nm led lamp (M455, Thorlabs) was used as a light source, which significantly improved the contrast of the images due to the large absorption band of the gel centered at 440 nm [6]. Samples, initially in the liquid state, were loaded into custom optical cells consisting of a microscope borosilicate glass slide and a cover slip separated by a 200 µm thick double-sided adhesive spacer to ensure airtightness (Gene Frame, Thermo Scientific). The sol-gel transition was induced by thermally quenching the sample using a dedicated Peltier cell (MHCS120i, Microptik). Image analysis was conducted using MATLAB (version: 9.13.0.2320565) and the Image Processing Toolbox (version 11.6). Images were segmented by first removing large background structures using a bottom-hat filter. Second, the branched structure of the growing spherulites was ”filled” using successive morphological erosion and dilatation steps, which afterwards improved their contour detection. Finally, thresholding was performed using an hysteresis method. The edge detection of the spherulites in the binary images was then performed using the matlab function regionprops, allowing to calculate their area $A$ as function of time and an equivalent sphere radius $r=\sqrt{A/\pi}$. The matlab code used for the image analysis can be found in Supplementary materials.
@@ II.5 Small Angle X-ray Scattering (SAXS)
Small Angle X-ray Scattering was conducted at the SWING beamline (Synchrotron Soleil, Saint-Aubin, France) [51]. The samples were enclosed in 1 mm borosilicate capillaries (WJM-Glass Muller GmbH) and positioned in a thermally controlled holder (Linkam Scientific Instruments, Tadworth, UK). The beam energy was set to 12 KeV. To derive the scattering intensity $I(q)$ as a function of norm of the scattering wave vector $q$, the two-dimensional scattering profiles of the viologen suspension and the background solvent were subtracted. The resulting scattering intensity remained isotropic, and an azimuthal average was performed to obtain a one-dimensional $I(q)$.
@@ III.4 Gelation kinetics
Figure 4: Gelation kinetics assessed by SAXS upon cooling at 0.5C///s. The azimuthally averaged intensity $I(q)$ is displayed vs. the wave vector $q$ at different times during the cooling ramp. The first scattered signal was measured at T = 27C (bottom curve) and acquisitions were performed every $5\leavevmode\nobreak\ \rm s$. Indicated time corresponds to the time elapsed since the beginning of the temperature quench from $T=90\leavevmode\nobreak\ \rm^{\circ}C$. Red curves represent the form factor of a core shell cylinder with constant parameters: core radius $r_{i}=5.5\leavevmode\nobreak\ \rm$, shell thickness $t_{s}=2.9\leavevmode\nobreak\ \rm nm$ and cylinder length $L=700\leavevmode\nobreak\ \rm nm$.
Hallow fibers nucleation – Using time-resolved SAXS during the temperature quench, we monitor the formation of the hallow fibers, Fig 4. The analysis of the scattering intensity $I(q)$ indicates that the gelators assemble directly into hollow tubes. Their geometrical properties, the core radius $r_{i}=5.5\leavevmode\nobreak\ \rm$ and the shell thickness $t_{s}=2.9\leavevmode\nobreak\ \rm nm$ remains unchanged during the kinetics. This evidences that crystalline fibers self-assemble into hollow cylinders, without intermediary structure allowing us to rule out mechanisms where the gelators self-assembles into large 2D sheets that subsequently fold into tubes [57, 58]. We also observe that the intensity of the scattering intensity increases with time indicating that the fibers concentration increases to rapidly form spherulite on larger length scales.
@@
Figure 5: Growth kinetics of spherulites assessed by bright field microscopy measurements. (a) Example of edge detection during the growth of a spherulite at $T=45^{\circ}$. From top to bottom, $t_{1}<t_{2}<t_{3}$. (b) Equivalent radius $r$ of the spherulite displayed in (a) as function of time. The times corresponding to each image of (a) are indicated in magenta. Plain black curve is the best linear fit of the data obtained from the tracking algorithms. Dotted black curves is an extrapolation of the linear growth of the spherulite up to its final size. (c) Nucleation time $t_{n}$ as function of the temperature of quench. (d) Growth rate $\overline{v}$ of individual spherulites as function of the temperature of quench. (e) Final radius $\overline{r}_{\infty}$ of the spherulites as function of the temperature of quench. Error bars are standard deviations.
@@
Those variations were also observed in other systems [38, 39, 40] and can be rationalized with variations of the driving force $\sigma$ upon changing the quench temperature. When the growth process is controlled by surface kinetics, it was concluded that variations in the driving force have much more effects on the nucleation rate $J$ than on the growth speed [40]. This is consistent with the small effect of $T$ on $\overline{v}$ and the increase of $t^{micro}_{N}$ with $T$. Associated with the lower nucleation rate at high $T$, the spherulites have less neighbors and are allowed to grow further, explaining the dependence of $\overline{r_{\infty}}$ with $T$. Besides, SALS experiments have shown that the fractal dimension of the spherulites decreases with $T$. For an heterogeneous nucleation mechanisms, the average branching length $l$ depends on the speed growth and on the nucleation rate according $l=\overline{v}/J$. Consequently, as $\overline{v}$ is roughly constant and $J$ decreases with $T$, the spherulite formed at high quench temperature are less branched with a lower fractal dimension.
@@
Figure 6: Growth kinetics assessed by UV-vis absorption spectroscopy measurements. (a) Absorbance spectra recorded over time after thermal quench at 35C. (b) Crystallinity of the system $X_{abs}$ as a function of the time elapsed since the nucleation time $t_{n}$. Colors code for the temperature of quench: T = 35 (yellow), 40 (orange) 45 (red) and 50 C (violet). $X_{abs}$ is calculated from the time-evolution of the absorbance of the charge transfer band at $\lambda=440$ nm, according to $X_{abs}(t)=[A_{CT}(t)-A_{CT}(t=0)]/[A_{CT}(t=\infty)-A_{CT}(t=0)]$, with $A_{CT}(t=0)$ and $A_{CT}(t=\infty)$ the absorbance at $t=t_{N}^{abs}$ and $t=\infty$. Black curves represent the Avrami relation which parameters are determined in (c)-(f). Inset displays a semi-log plot of the same data. (c)-(f) Avrami plots of the data displayed in (b). Black line is the best power law fit of the data. The power exponent $p_{abs}$ is displayed in (g) as function of the quench temperature $T$. (h) Nucleation time $t_{N}^{abs}$ as function of the temperature of quench.
@@
Inspired by earlier reports [2, 47], we use the Avrami representation $-\ln(1-X_{abs})$ vs. $t-t^{abs}_{N}$ [Fig. 6(c)-(f)]. In this representation, we observe three regimes. At short time scales, $-\ln(1-X_{abs})$ evolves slowly with time. At intermediate time scales $-\ln(1-X_{abs})$ follows a power law evolution with an exponent $p_{abs}$: $\ln(1-X)=k(t-t^{abs}_{N})^{p_{abs}}$. At long time scales $-\ln(1-X_{abs})$ tends to plateau as the gelation reaction is finished. The early time scales regime pertains to the initial stage of the self-assembly of the gelator into hollow tubes. We propose that a cooperative supramolecular polymerization mechanism occurs during this process, where the formation of a tubular structure strongly promotes the self-assembly of $Chol_{2}V^{2+}$[61], resulting in an apparent ’instantaneous growth’ of the fibers. This hypotheses is backed up by two observations. First, when we plot $A(t)-A(t=0)$ (see Appendix V.4), thus removing the contribution of the peak at $\lambda=370$ nm, we observe that both the shape and position along the $\lambda$-axis of the broad adsorption peak at $\lambda=440$ nm remain unchanged. This suggests that at the interaction scales, the viologen gelators consistently interact in the same manner and likely do not assemble into structures other than hollow tubes in agreement with the time resolved SAXS presented in Fig. 4. The intermediate time scale regime allows us to determine the Avrami exponent $n$ in Eq. 2, here denoted $p_{abs}$. In Fig. 6(g), $p_{asb}$ is consistent with the fractal dimension $d=p_{SALS}$ determined from scattering measurements, and decreases with $T$ [Fig. 2(e)]. Given that $n_{G}=1$ as determined from microscopy measurements, and since we identified that $d=p_{abs}=p_{SALS}$, this suggests that $n_{N}\approx 0$, indicating either pre-existing nuclei conditions or a decrease in the nucleation rate.
@@
Evolution of the gel mechanical properties – Finally, we assess the macroscopic kinetics of gelation using rheometric measurements. Fig. 7(a) displays the temporal evolution of the elastic modulus $G^{\prime}$ and the viscous modulus $G^{\prime\prime}$ for different quench temperatures. At short time scales, $G^{\prime\prime}$ dominates over $G^{\prime}$, while at longer time scales, $G^{\prime}$ surpasses $G^{\prime\prime}$. This behavior is characteristic of a sol-gel transition, and we estimate the gel time $t_{g}$ as the time for which $G^{\prime}>G^{\prime\prime}$. The gelation times displayed in Fig. 7(h) are of the same order of magnitude than the nucleation times determined from microscopy and UV-Vis absorption spectroscopy measurements, as expected from the large growth speed of the fibers ($\approx 2$ µm.s-1). At low quench temperatures, the viscoelastic moduli undergoes a small decrease after reaching a maximum value [Fig. 7(a)]. This behavior is attributed to negative normal forces, resulting from a contraction of the gel as reported for bio-polymer gels [62]. In the inset graph of Fig. 7(a), we rescale the elastic modulus by the absorbance value $A_{CT}$ to account for variations in the concentration of crystallized gelator with $T$. When displayed against the rescaled time $t/t_{g}$, $\tilde{G^{\prime}}$ reaches a plateau value equals to about 1000 Pa for all quench temperatures. In previous reports, the elasticity of multi-domain networks increased with a decrease of $\overline{r}_{\infty}$ [5]. Given the strong variations of the final spherulite radius $\overline{r}_{\infty}$ and fractal dimension with $T$ [Fig. 5(e) and Fig. 2(e)], the present results suggest that the mesoscopic structure of the gel has little effect of its mechanical properties. However, potential differences may be masked by the small value of the gap ($e=500$ µm). The final spherulite radius ($200\leq\overline{r}_{\infty}\leq 800$ µm) approximately matches the value of $e$, indicating that rheometric measurements predominantly assess the elasticity of the fibers or more localized interactions, such as fiber side branching, rather than the elasticity of the spherulites themselves [34, 5, 43, 63]. In line with the work of Liu and coworkers [2, 3, 64], the fraction of crystalline phase $X$ of a fiber network could be deduced from the norm of the complex viscosity $\eta^{*}=(G^{\prime}+iG^{\prime\prime})/\omega$, where $\omega$ is the angular frequency of small amplitude oscillation used to measure the viscoelastic moduli. By applying Einstein’s relation ($\eta/\eta_{0}=(1+2.5\phi)$), one may connect the normalized viscosity and volume fraction of spherulite $\phi$. Accordingly, the fraction of crystalline phase at a given time is calculated as $X_{rheo}=\frac{|\eta^{*}(t)|-|\eta^{*}(t=0)|}{|\eta^{*}(t=\infty)|-|\eta^{*}(t=0)|}$. In Fig. 7(b), $X_{rheo}$ is displayed as a function of the reduced time $t-t_{g}$ for different quench temperatures. As with UV-vis absorption spectroscopy measurements, we use Avrami plots $-\ln(1-X_{rheo})$ vs $t-t_{g}$ to determine the Avrami exponents [Fig 7(c)-(e)]. In Fig. 7(g), we found that the exponents determined from rheological measurement $p_{rheo}$ are comprised between 3.6 and 2.3, and are thus overestimated in comparison with exponents determined from scattering and UV-Vis absorption spectroscopy measurements (Fig. 2(d-inset) and Fig. 6(g)). We cannot identify $p_{rheo}$ to $d$ as we identified $p_{abs}$ and $p_{SALS}$ to $d$ in the Avrami equation Eq. 2. We believe that these discrepancies stem from the assumption of the linear relation between the viscosity of the solution and the volume fraction $\phi$, which is not valid in the present system. Indeed, the Einstein’s relation was derived in the case of a hard sphere suspension in diluted regime. While the spherulites may behave like a dilute suspension during the early stages of crystallization (i.e., $t<<t_{g}$), the spherulites grow and interact to form network for $t\geq t_{g}$, with numerous relaxation modes [65].

### v2

@@
Supramolecular crystal gels, a subset of molecular gels, form through self-assembly of low molecular weight gelators into interconnecting crystalline fibers, creating a three-dimensional soft solid network. This study focuses on the formation and properties of viologen-based supramolecular crystalline gels. It aims to answer key questions about the tunability of network properties and the origin of these properties through in-depth analyses of the gelation kinetics triggered by thermal quenching. Experimental investigations, including UV-Vis absorption spectroscopy, rheology, microscopy and scattering measurements, contribute to a comprehensive and self-consistent understanding of the system kinetics. We confirm that the viologen-based gelators crystallize by forming nanometer radius hollow tube that assemble into micro to millimetric spherulites. We then show that the crystallization follows the Avrami theory and is based on pre-existing nuclei. We also establish that the growth is interface controlled leading to the hollow tubes to branch into spherulites with fractal structures. Finally, we demonstrate that the gel properties can be tuned depending on the quenching temperature. Lowering the temperature results in the formation of denser and smaller spherulites. In contrast, the gels elasticity is not significantly affected by the quench temperature, leading us to hypothesize that the spherulites densification occurs at the expense of the connectivity between spherulite
@@ I Introduction
[…]
@@
However, it remains unclear how these spherulites contribute to the formation of a 3D network. Do the branches entangle between neighboring spherulites, or do branches stick to one another due to attractive interactions? Predictions [29, 21, 22, 37] were made at the macroscopic level to establish a direct link between the crystal gel structure and its mechanical properties. In this context, the spherulites are best described as Cayley trees of size $\xi$ and fractal dimension $d_{f}$ composed of branching segment of length $l$ and radius $a$. Based on simulations, Shi et al. [46] assimilated the gel elasticity to that of the spherulites. In [29, 21, 22, 37], the authors proposed a model derived from the Avrami model [3] that links the temporal evolution of the elastic modulus to the fractal dimension $d_{f}$ of the crystal gel network. The model establishes that $X$, the ratio of the volume fraction of crystal material at time $t$ to the volume fraction at $t\to\infty$, is linked to the evolution of the norm of the complex viscosity $\eta^{*}$ through the equations $X=\frac{|\eta^{*}|-|\eta^{*}(0)|}{|\eta^{*}_{\infty}|-|\eta^{*}(0)|}$ and $-\ln(1-X)=k(t-t_{g})^{d_{f}}$, assuming $n_{N}=0$ and $n_{G}=1$ et Eq. 2.
@@ II.1 Viologen-based gelator
The viologen-based gelator used for all experiments is a cholesteryl-substituted viologen noted $Chol_{2}V^{2+}$ (Fig. 1a). It was obtained as a bromide salt $V(Br)_{2}$ using a protocol described previously  [2]. The gelation ability of $Chol_{2}V^{2+}$ was first reported by Xue et al., who demonstrated that the gel fibrillar network could be efficiently used as a template for obtaining monodisperse mesoporous silica [59]. Our previous work showed that self-assembly of $Chol_{2}V^{2+}$ in 1-Pentanol relies on the formation of charge transfer (CT) complexes between $Chol_{2}V^{2+}$ (acceptor unit) and halide ions, $Br^{-}$ or $I^{-}$ (donor unit). The anions associated with the dicationic species $Chol_{2}V^{2+}$ play a major role in the self-assembly process and that the rheological, optical and electrical properties of the gels can be tuned by the addition of salts. We also showed that the addition of lithium bis(trifluoromethanesulfonyl) ($LiTFSI$) helps to increase the solubility of $Chol_{2}V^{2+}$ in pentanol but that too much TFSI anions can prevent gelation. Addition of $TBAI$ also increases the amount of donors species available to form the $[Chol_{2}V^{2+}/I^{-}]$ CT complexes promoting gelation but yields opaque gels displaying high gelation temperature. An optimal composition of $TBAI$ and $LiTFSI$ has thus been determined to obtain a translucent and robust gel featuring a moderate gelation temperature and a high ionic conductivity [2]. This study specifically focuses on this optimal composition consisting of $Chol_{2}V(Br)_{2}$ (1.3 mM, 0.2 wt%), $TBAI$ (50 mM), and $LiTFSI$ (100 mM) in 1-Pentanol. In all experiments, the sol-gel transition was induced by thermally quenching homogeneous solutions at a rate of $\sim 0.6^{\circ}$/s from 80 C to a final temperature $T$ using a Peltier module. The shallower quench where carried out to a final temperature $T=50^{\circ}$ while the deeper quenches where carried out down to $T=20^{\circ}$. The time $t=0$ s is defined as the moment when the final temperature is reached.
@@ II.4 Bright field microscopy
Microscopy measurements were performed using a Nikon Ti-eclipse inverted light microscope operating in bright-field mode. A Zyla camera from Andor, equipped with a 10× objective, was employed to capture images of the samples. A 455 nm led lamp (M455, Thorlabs) was used as a light source, which significantly improved the contrast of the images due to the large absorption band of the gel centered at 440 nm [2]. Samples, initially in the liquid state, were loaded into custom optical cells consisting of a microscope borosilicate glass slide and a cover slip separated by a 200 µm thick double-sided adhesive spacer to ensure airtightness (Gene Frame, Thermo Scientific). The sol-gel transition was induced by thermally quenching the sample using a dedicated Peltier cell (MHCS120i, Microptik). Image analysis was conducted using MATLAB (version: 9.13.0.2320565) and the Image Processing Toolbox (version 11.6). Images were segmented by first removing large background structures using a bottom-hat filter. Second, the branched structure of the growing spherulites was ”filled” using successive morphological erosion and dilatation steps, which afterwards improved their contour detection. Finally, thresholding was performed using an hysteresis method. The edge detection of the spherulites in the binary images was then performed using the matlab function regionprops, allowing to calculate their area $A$ as function of time and an equivalent sphere radius $r=\sqrt{A/\pi}$. We define $t^{mciro}_{N}$ as the earliest time at which a spherulite can be consistently detected. $t^{mciro}_{N}$ is influenced by the tracking algorithm’s properties: its spatial sensitivity around r=25 $\mu$m and its ability to track the growth of a single spherulite over at least 20 frames.
@@ II.5 Small Angle X-ray Scattering (SAXS)
Small Angle X-ray Scattering was conducted at the SWING beamline (Synchrotron Soleil, Saint-Aubin, France) [54]. The samples were enclosed in 1 mm borosilicate capillaries (WJM-Glass Muller GmbH) and positioned in a thermally controlled holder (Linkam Scientific Instruments, Tadworth, UK). The beam energy was set to 12 KeV. The camera (EigerX4M with a detection area: 162.5 $\times$ 155.2 mm2 and a pixel size : 75 $\times$ 75 $\mu$m2) was set at a distance of 1.5m from the sample. To derive the scattering intensity $I(q)$ as a function of norm of the scattering wave vector $q$, the two-dimensional scattering profiles of the viologen suspension and the background solvent were subtracted. The resulting scattering intensity remained isotropic, and an azimuthal average was performed to obtain a one-dimensional $I(q)$.
@@ III.4 Gelation kinetics
Figure 4: Gelation kinetics assessed by SAXS upon cooling at 0.5C///s. The azimuthally averaged intensity $I(q)$ is displayed vs. the wave vector $q$ at different times during the cooling ramp to 20C. The time $t_{0}=140$ s corresponds to the time necessary to reach 20C from 90C. The first scattered signal was measured 15 s before reaching the quench temperature at T = 27C (bottom curve). The red curve represents the fit of the form factor for a core-shell cylinder with a core radius $r_{i}=5.5\leavevmode\nobreak\ \rm$, a shell thickness $t_{s}=2.9\leavevmode\nobreak\ \rm nm$ and a cylinder length $L>700\leavevmode\nobreak\ \rm nm$ for $t-t_{0}=10$ s. This fit with same parameters remains valid at all time scales.
Hollow fibers nucleation – Using time-resolved SAXS during the temperature quench, we monitor the formation of the hollow fibers, Fig 4. The analysis of the scattering intensity $I(q)$ indicates that the gelators assemble directly into hollow tubes. Their geometrical properties, the core radius $r_{i}=5.5\leavevmode\nobreak\ \rm$ and the shell thickness $t_{s}=2.9\leavevmode\nobreak\ \rm nm$ remains unchanged during the kinetics. This evidences that crystalline fibers self-assemble into hollow cylinders, without intermediary structure allowing us to rule out mechanisms where the gelators self-assembles into large 2D sheets that subsequently fold into tubes [20, 23]. We also observe that the intensity of the scattering intensity increases with time indicating that the fibers concentration increases to rapidly form spherulite on larger length scales.
@@
Figure 5: Growth kinetics of spherulites assessed by bright field microscopy measurements. (a) Example of edge detection during the growth of a spherulite at $T=45^{\circ}$. From top to bottom, $t_{1}<t_{2}<t_{3}$. (b) Equivalent radius $r$ of the spherulite displayed in (a) as function of time. The times corresponding to each image of (a) are indicated in magenta. Plain black curve is the best linear fit of the data obtained from the tracking algorithms. Dotted black curves is an extrapolation of the linear growth of the spherulite up to its final size. (c) Nucleation time $t^{mciro}_{N}$ as function of the temperature of quench. (d) Growth rate $\overline{v}$ of individual spherulites as function of the temperature of quench. (e) Final radius $\overline{r}_{\infty}$ of the spherulites as function of the temperature of quench. Error bars are standard deviations.
@@
Those variations were also observed in other systems [32, 53, 15] and can be rationalized with variations of the driving force $\sigma$ upon changing the quench temperature. When the growth process is controlled by surface kinetics, it was concluded that variations in the driving force have much more effects on the nucleation frequency $J$ [s-1] than on the growth speed [15]. This is consistent with the small effect of $T$ on $\overline{v}$ and the increase of $t^{micro}_{N}$ with $T$. Associated with the lower nucleation frequency at high $T$, the spherulites have less neighbors and are allowed to grow further, explaining the dependence of $\overline{r_{\infty}}$ with $T$. Besides, SALS experiments have shown that the fractal dimension of the spherulites decreases with $T$. For an heterogeneous nucleation mechanisms, the average branching length $l$ depends on the speed growth and on the nucleation frequency according $l=\overline{v}/J$. Consequently, as $\overline{v}$ is roughly constant and $J$ decreases with $T$, the spherulite formed at high quench temperature are less branched with a lower fractal dimension.
@@
Figure 6: Growth kinetics assessed by UV-vis absorption spectroscopy measurements. (a) Absorbance spectra recorded over time after thermal quench at 35C. The color codes for the time $t$. $t=0$ s correspond to the time at which $T=35^{\circ}$. (b) Crystallinity of the system $X_{abs}$ as a function of the time elapsed since the nucleation time $t_{n}$. Colors code for the temperature of quench: T = 35 (yellow), 40 (orange) 45 (red) and 50 C (violet). $X_{abs}$ is calculated from the time-evolution of the absorbance of the charge transfer band at $\lambda=440$ nm, according to $X_{abs}(t)=[A_{CT}(t)-A_{CT}(t=0)]/[A_{CT}(t=\infty)-A_{CT}(t=0)]$, with $A_{CT}(t=0)$ and $A_{CT}(t=\infty)$ the absorbance at $t=t_{N}^{abs}$ and $t=\infty$. Black curves represent the Avrami relation which parameters are determined in (c)-(f). Inset displays a semi-log plot of the same data. (c)-(f) Avrami plots of the data displayed in (b). Black line is the best power law fit of the data. The power exponent $p_{abs}$ is displayed in (g) as function of the quench temperature $T$. (h) Nucleation time $t_{N}^{abs}$ as function of the temperature of quench.
@@
Inspired by earlier reports [27, 21], we use the Avrami representation $-\ln(1-X_{abs})$ vs. $t-t^{abs}_{N}$ [Fig. 6(c)-(f)]. In this representation, we observe three regimes. At short time scales, $-\ln(1-X_{abs})$ evolves slowly with time. At intermediate time scales $-\ln(1-X_{abs})$ follows a power law evolution with an exponent $p_{abs}$: $\ln(1-X)=k(t-t^{abs}_{N})^{p_{abs}}$. At long time scales $-\ln(1-X_{abs})$ tends to plateau as the gelation reaction is finished. The early time scales regime pertains to the initial stage of the self-assembly of the gelator into hollow tubes. We propose that a cooperative supramolecular polymerization mechanism occurs during this process, where the formation of a tubular structure strongly promotes the self-assembly of $Chol_{2}V^{2+}$[9], resulting in an apparent ’instantaneous growth’ of the fibers. This hypotheses is backed up by two observations. First, when we plot $A(t)-A(t=0)$ (see Appendix V.4), thus removing the contribution of the peak at $\lambda=370$ nm, we observe that both the shape and position along the $\lambda$-axis of the broad adsorption peak at $\lambda=440$ nm remain unchanged. This suggests that at the interaction scales, the viologen gelators consistently interact in the same manner and likely do not assemble into structures other than hollow tubes in agreement with the time resolved SAXS presented in Fig. 4. The intermediate time scale regime allows us to determine the Avrami exponent $n$ in Eq. 2, here denoted $p_{abs}$. In Fig. 6(g), $p_{asb}$ is consistent with the fractal dimension $d=p_{SALS}$ determined from scattering measurements, and decreases with $T$ [Fig. 2(e)]. Given that $n_{G}=1$ as determined from microscopy measurements, and since we identified that $d=p_{abs}=p_{SALS}$, this suggests that $n_{N}\approx 0$, indicating either pre-existing nuclei conditions formed during the early time scales or a decrease in the nucleation frequency.
@@
Evolution of the gel mechanical properties – Finally, we assess the macroscopic kinetics of gelation using rheometric measurements. Fig. 7(a) displays the temporal evolution of the elastic modulus $G^{\prime}$ and the viscous modulus $G^{\prime\prime}$ for different quench temperatures. At short time scales, $G^{\prime\prime}$ dominates over $G^{\prime}$, while at longer time scales, $G^{\prime}$ surpasses $G^{\prime\prime}$. This behavior is characteristic of a sol-gel transition, and we estimate the gel time $t_{g}$ as the time for which $G^{\prime}>G^{\prime\prime}$. The gelation times displayed in Fig. 7(h) are of the same order of magnitude than the nucleation times determined from microscopy and UV-Vis absorption spectroscopy measurements, as expected from the large growth speed of the fibers ($\approx 2$ µm.s-1). At low quench temperatures, the viscoelastic moduli undergoes a small decrease after reaching a maximum value [Fig. 7(a)]. This behavior is attributed to negative normal forces, resulting from a contraction of the gel as reported for bio-polymer gels [33]. In the inset graph of Fig. 7(a), we rescale the elastic modulus by the absorbance value $A_{CT}$ to account for variations in the concentration of crystallized gelator with $T$. When displayed against the rescaled time $t/t_{g}$, $\tilde{G^{\prime}}$ reaches a plateau value equals to about 1000 Pa for all quench temperatures. In previous reports, the elasticity of multi-domain networks increased with a decrease of $\overline{r}_{\infty}$ [26]. Given the strong variations of the final spherulite radius $\overline{r}_{\infty}$ and fractal dimension with $T$ [Fig. 5(e) and Fig. 2(e)], the present results suggest that the mesoscopic structure of the gel has little effect of its mechanical properties. However, potential differences may be masked by the small value of the gap ($e=500$ µm). The final spherulite radius ($200\leq\overline{r}_{\infty}\leq 800$ µm) approximately matches the value of $e$, indicating that rheometric measurements predominantly assess the elasticity of the fibers or more localized interactions, such as fiber side branching, rather than the elasticity of the spherulites themselves [61, 26, 46, 62]. In line with the work of Liu and coworkers [27, 28, 30], the fraction of crystalline phase $X$ of a fiber network could be deduced from the norm of the complex viscosity $\eta^{*}=(G^{\prime}+iG^{\prime\prime})/\omega$, where $\omega$ is the angular frequency of small amplitude oscillation used to measure the viscoelastic moduli. By applying Einstein’s relation ($\eta/\eta_{0}=(1+2.5\phi)$), one may connect the normalized viscosity and volume fraction of spherulite $\phi$. Accordingly, the fraction of crystalline phase at a given time is calculated as $X_{rheo}=\frac{|\eta^{*}(t)|-|\eta^{*}(t=0)|}{|\eta^{*}(t=\infty)|-|\eta^{*}(t=0)|}$. In Fig. 7(b), $X_{rheo}$ is displayed as a function of the reduced time $t-t_{g}$ for different quench temperatures. As with UV-vis absorption spectroscopy measurements, we use Avrami plots $-\ln(1-X_{rheo})$ vs $t-t_{g}$ to determine the Avrami exponents [Fig 7(c)-(e)]. In Fig. 7(g), we found that the exponents determined from rheological measurement $p_{rheo}$ are comprised between 3.6 and 2.3, and are thus overestimated in comparison with exponents determined from scattering and UV-Vis absorption spectroscopy measurements (Fig. 2(d-inset) and Fig. 6(g)). We cannot identify $p_{rheo}$ to $d$ as we identified $p_{abs}$ and $p_{SALS}$ to $d$ in the Avrami equation Eq. 2. We believe that these discrepancies stem from two assumptions. First, $t_{g}$ might not be the right characteristic time to normalise the $x$-axis in the Avrami representation. Replacing $t_{g}$ by $t_{N}^{abs}$ does not improve the results; in fact, it makes them even worse. Since $t_{g}$ is mostly larger than $t_{N}^{abs}$, plotting $-ln(1-X_{rheo})$ as a function of $t-t_{N}^{abs}$ leads to even higher values of $p_{rheo}$, which is already too large. Second, the assumption of the linear relation between the viscosity of the solution and the volume fraction $\phi$, is not valid in the present system. Indeed, the Einstein’s relation was derived in the case of a hard sphere suspension in diluted regime. While the spherulites may behave like a dilute suspension during the early stages of crystallization (i.e., $t<<t_{g}$), the spherulites grow and interact to form network for $t\geq t_{g}$, with numerous relaxation modes [49].